OpenStudy (anonymous):

For the polynomial shown below, find f(-2) f(x)=3x^4-x^2+4x-2

OpenStudy (anonymous):

\[f(-2)=3(-2)^{4}-(-2)^{2}+4(-2)-2=3\times16-4-8-2=3\times16-14=48-14=34\]

OpenStudy (anonymous):

=48-14=34